An aluminum wing on a passenger jet is 30m long when its temperature is 20 degrees C. At what temperature would the wing be 5 cm (0.05) shorter?

Coefficient of Aluminium = 25 X10^-6
Answer #1

/\L = alpha * L * /\T

0.05 = 0.000025 * 30 * /\T

0.05 = 0.00075 * /\T

0.05 / 0.00075 = /\T

/\T = 66.67 ---> 20 - 66.67 = -46.67 C
Answer #2

//\L = alpha * L * /\T

0.05 = 0.000025 * 30 * /\T

0.05 = 0.00075 * /\T

0.05 / 0.00075 = /\T

/\T = 66.67 ---> 20 - 66.67 = -46.67 C


. because the temperature is below zero, and probably below you have ever encountered doesn't mean it doesn't exist. A temperature of -46is cold, but not extremely cold (theoretical lowest temperature possible -273.15 C).
A secondary reason is that the temperature should be lower than the start temperature.heat expands things